OpenStudy (anonymous):

I'm really confused 0_o (i suck horribly at Algebra) • Solve each inequality 1. w + 3 > 9 2. v - 6 < 4 3. -4 < t + 8 4. n - ½ ≥ ¾ 5. 22.3 ≤ 13.7 + h 6. q + 0.5 > -2 Thanks ♥ ☺

OpenStudy (anonymous):

w>3

OpenStudy (anonymous):

v<10

OpenStudy (anonymous):

-12<t

OpenStudy (anonymous):

8.6 is less then or equal to h

OpenStudy (anonymous):

q>-1.5

OpenStudy (anonymous):

thank you @Gorge :) ^_^ ♥

OpenStudy (anonymous):

No problem. That what I'm here for

OpenStudy (anonymous):

1. w + 3 > 9 w + 3 - 3 > 9 - 3 w > 6 2. v - 6 < 4 v - 6 + 6 < 4 + 6 v < 10 3. - 4 < t + 8 -8 - 4 < t + 8 - 8 - 12 < t 4. n - 1/2 <= 3/4 (4) n - 1/2 <= (4)(3/4) 4n - 2 <= 3 4n - 2 + 2 <= 3 + 2 4n <= 5 n < = 5/4 ( I think) 5. 22.3 <= 13.7 +h 22.3 - 13.7 <= 13.7 - 13.7 + h 8.6 <= h 6. q + .5 > -2 q + .5 - .5 > -2 - .5 q > - 2.5

OpenStudy (anonymous):

omg thanks ^_^ i appreciate it (i probably spelled that wrong 0_o)